# NeuronSeek: On Stability and Expressivity of Task-driven Neurons

### v1

@@ II-C Super-expressive Activation Functions
III Task-driven Neuron Based on Tensor Decomposition (NeuronSeek-TD)
III-A NeuronSeek-SR
In this section, we briefly illustrate how NeuronSeek-SR creates novel neurons. It follows a two-stage process, as illustrated in Figure 1. First, the vectorized SR identifies the optimal homogeneous neuronal formula by fitting task-specific data. Second, the derived formula is implemented as the neuron’s aggregation function with learnable coefficients.
III-A1 Vectorized SR
Conventional SR encodes a mathematical expression as a tree-structured directed graph and utilizes GP to navigate the solution space [15]. The GP optimization process employs biologically-inspired evolutionary operators, including: i) crossover operations that exchange subtrees between candidate solutions to maintain population diversity, and ii) mutation operations that introduce random modifications to explore novel regions of the function space. While this approach offers notable flexibility for symbolic optimization, it faces significant challenges such as computational complexity, formula heterogeneity, and overfitting when applied to high-dimensional problems. All these together are not ideal for the task of prototyping new neuronal formulations.
Unlike the conventional SR, which discovers heterogeneous formulas for each variable, the vectorized SR regularizes all input variables to learn the same formula [13, 14]. This single change brings huge advantages: i) Lower complexity. Homogeneous formulas generated by the vectorized SR have significantly fewer parameters, scaling linearly with the input dimensionality. ii) Faster computation and minor overfitting risks. By limiting the search space, the vectorized SR reduces the risk of overfitting and also makes the regression process faster and more computationally efficient, especially for high-dimensional data.
III-A2 Parameterization
The learned formulas from the vectorized SR are parameterized, i.e., making the coefficients trainable. This allows neurons to fine-tune their behavior in the context of a network. While task-based neurons alone may have limited approximation power, connecting them into a neural network framework will further amplify the power of neurons, thereby enabling efficient optimization and improving performance in handling complex tasks.
Despite the success of the vectorized SR, its built-in tree-based evolutionary algorithm still suffers from poor convergence in high-dimensional space. Consequently, the derived formulas exhibit significant sensitivity to key hyperparameters—including initialization conditions, population size, and mutation rate—often producing inconsistent solutions across different runs. This inherent instability in the NeuronSeek-SR framework stems from the intrinsic issues of genetic algorithms.
III-B NeuronSeek-TD
In this work, we propose NeuronSeek-TD, a faster and more stable alternative to conventional vectorized SR for neuronal formula discovery. We reasonably assert that the maximum required polynomial order for any given task is bounded by a constant $N$. By applying tensor decomposition to optimize coefficients, we reformulate the unstable formula search problem into a stable low-rank coefficient optimization task. Furthermore, we augment the polynomial structure through elementary functions, providing greater design flexibility and improved performance in high-dimensional pattern recognition tasks. The overall framework is illustrated in Figure 2, which comprises two stages: search for the neuronal formula and construct the task-driven neural network.
III-B1 Neuronal formula discovery
In the first stage, we adopt a polynomial model to identify the expressions that best fit the task-specific data. For consistent processing across modalities, both tabular and image inputs are flattened into a vector representation $\mathbf{z}\in\mathbb{R}^{d}$. Then, given the input data $\mathbf{z}$, the initial neuronal formula is defined as follows:
$f(\mathbf{z})$ $=C_{1}\cdot\beta_{j}+C_{2}\cdot\mathbf{z}^{\top}\mathbf{W}^{[1]}+C_{3}\cdot\mathbf{z}^{\top}\mathbf{W}^{[2]}\mathbf{z}$ (2)
$+C_{4}\cdot\mathbf{W}^{[3]}\times_{1}\mathbf{z}\times_{2}\mathbf{z}\times_{3}\mathbf{z}+\cdots+C_{N}\cdot\mathbf{W}^{[N]}\prod_{n=1}^{N}\times_{n}\mathbf{z}$
$+D\cdot\sin(\mathbf{z}),$
where $\mathbf{W}^{[n]}$ are weights and $C_{j}$ are coefficients. To put it in a more compact way, we have
$f(\mathbf{z})$ $=\sum_{n=1}^{N}\left(C_{n}\cdot\mathcal{W}^{[n]}\prod_{m=1}^{n}\mathbf{z}\right)+D\cdot\sin(\mathbf{z}),$ (3)
where the order $N$ and the rank are two hyperparameters to be set appropriately. The main idea of this method is to outer-product a series of rank-1 tensors to represent high-order tensors so that the computation cost is reduced significantly. The tensor decomposition process could be formulated as follows:
$\mathbf{W}^{[n]}=\sum_{r=1}^{R}\mathbf{a}_{r}^{(1)}\circ\mathbf{a}_{r}^{(2)}\circ\dots\circ\mathbf{a}_{r}^{(n)},$ (4)
where $R$ is decomposition rank, and $n$ is the order of the decomposed tensor $\mathbf{W}$. Here, we use COMP/PARAFAC (CP) decomposition to execute this process. CP decomposition has been widely adopted for modeling multilinear relationships in machine learning applications [51, 31]. In contrast to Tucker decomposition, CP decomposition provides both computational and structural advantages for our neuronal formula discovery framework. Computationally, CP decomposition requires only $O(nRd)$ parameters for an $n$-th order tensor of size $d^{n}$, compared to Tucker’s $O(R^{n}+nRd)$ complexity, enabling efficient handling of high-order polynomial terms. An example of CP decomposition is illustrated in Figure 4. While other tensor decomposition approaches could be explored (e.g., tensor ring decomposition, hierarchical Tucker decomposition) [52, 53], the efficiency and structural benefits of CP decomposition have already aligned well with the requirements of our polynomial modeling framework.
@@ III-B1 Neuronal formula discovery
Moreover, the incorporation of the trigonometric term $D\sin(\mathbf{z})$ serves a pivotal purpose. It extends the model’s capacity to capture periodic patterns and high-frequency components that intrinsically appear in many real-world datasets [54]. This is particularly crucial as pure polynomial expansions exhibit limited spectral sensitivity.
We then impose sparsity constraints through $L_{1}$ regularization on both the coefficients $C_{i}$ and weight tensors $\mathbf{W}^{[i]}$ to promote computational efficiency. Specifically, since each tensor $\mathbf{W}^{[i]}$ is decomposed using CP decomposition as $\mathbf{W}^{[i]}=\sum_{r=1}^{R}\mathbf{a}_{r}^{(1)}\circ\mathbf{a}_{r}^{(2)}\circ\dots\circ\mathbf{a}_{r}^{(i)}$, we apply sparsity regularization directly to the factor matrices. Thus, the complete optimization objective becomes
$\underset{\{\mathbf{a}_{r}^{(j)}\},C_{i}}{\min}\sum_{k=1}^{K}\mathcal{L}(f(\mathbf{x}_{k}),y_{k})+\lambda_{1}\|\mathbf{C}\|_{1}+\lambda_{2}\sum_{i=1}^{N}\sum_{r=1}^{R}\sum_{j=1}^{i}\|\mathbf{a}_{r}^{(j)}\|_{1},$ (5)
where $\mathcal{L}(\cdot)$ is the task-specific loss function, set to mean squared error (MSE) for regression and cross-entropy (CE) loss for classification. This structured sparsity approach encourages the discovery of interpretable and simplified polynomial formulas while maintaining computational efficiency through the low-rank tensor representation.
To further enhance the compactness of the discovered neuronal formula beyond the $L_{1}$ regularization’s sparsity induction, we introduce a post-training thresholding mechanism for explicit term selection. While the $L_{1}$ regularization effectively drives many coefficients toward zero, empirical observations reveal that a non-negligible portion of coefficients $C_{i}$ retain small but non-zero values, which cumulatively increase model complexity without substantial predictive contribution. The thresholding mechanism is a statistically grounded pruning strategy: After model convergence, we compute the mean $\mu_{|C|}$ and the standard deviation $\sigma_{|C|}$ of the absolute values of all coefficients $\{|C_{i}|\}_{i=1}^{N}$, then construct the final neuronal formula by retaining only terms whose coefficients satisfy $|C_{i}|\geq\tau$, where $\tau=\mu_{|C|}-\sigma_{|C|}$. This can further increase the stability of identified formulas.
Remark 1. Three key features of the proposed method warrant emphasis. i) The higher-order polynomial architecture systematically enables the learning of the polynomials. It is noticed that Eq. (2) is rather inclusive, e.g., it identifies multi-order variable interactions. This structural flexibility facilitates sophisticated nonlinear approximations, aligning the model more closely with the underlying data distribution. ii) The integration of trainable coefficients $\mathbf{C}$ with the $L_{1}$ regularization imposes selective sparsity constraints on the polynomial expansion. Specifically, the regularization term $\mathbf{C}$ induces sparsity by driving coefficients of less predictive terms toward zero, thereby automatically pruning nonessential higher-order components and yielding a parsimonious architecture. iii) The complementary $L_{1}$ regularization on the coefficients $\mathbf{W}$ enforces structured sparsity in the parameter space. This dual regularization strategy enhances feature selection while simultaneously improving computational efficiency—a critical advantage in high-dimensional contexts. Collectively, these design choices balance expressive capacity with compactness, ensuring robust performance without overparameterization.
III-B2 Task-driven network construction
In the second stage, the discovered neuronal formula is first parameterized into neuron-like structures. The selected polynomial and trigonometric terms are assigned trainable weights and integrated into neural network architectures. For tabular data processing, we employ fully connected neural networks. A polynomial with the maximum polynomial order of 3 is expressed as
${y}=\sigma(\mathbf{W}_{1}^{\top}\mathbf{z}+\mathbf{W}_{2}^{\top}\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{3}^{\top}\mathbf{z}\odot\mathbf{z}\odot\mathbf{z}+{b}),$ (6)
with $\sigma$ being the ReLU activation function. If a specific order term is absent in the neuronal formula, its corresponding weight and bias parameters are initialized to zero.
Moreover, we extend our framework to convolutional neurons for images. As mentioned, the trigonometric term is incorporated to model periodic patterns and high-frequency components inherent in high-dimensional images. Consequently, the complete neuronal formula is structured as follows:
${y}=\sigma(\mathbf{W}_{1}^{\top}*\mathbf{z}+\mathbf{W}_{2}^{\top}*\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{3}^{\top}*\mathbf{z}\odot\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{4}*\sin(\mathbf{z})+{b}),$ (7)
where * denotes the convolution operation.
Finally, we introduce a simple initialization strategy to enhance training stability. Followed by [33], the weight matrix $\mathbf{W}_{1}$ is initialized as $\mathbf{W}_{1}\sim\mathcal{N}(0,\sigma^{2})$, where $\sigma=\sqrt{1/32(k_{n})}$, while all other weights are initialized to a tiny number such as 0.001. This strategy ensures that the network initially behaves like linear neurons and gradually learns higher-order terms during training. This strategy has been demonstrated the effectiveness in training quadratic neurons [55].
IV Super-expressive Property Theory of Task-driven Neuron
@@
Step 1. As Figure 4 shows, in equal distance, divide $[0,1)$ into small intervals $\mathcal{J}_{k}=\left[\frac{k-1}{K},\frac{k}{K}\right)$ for $k\in\{1,2,\cdots,K\}$, where $K$ is the number of intervals. The left endpoint of $\mathcal{J}_{k}$ is $x_{k}$. A higher $K$ leads to a smaller approximation error. We construct a piecewise constant function $h$ to approximate $f$. The error can be arbitrarily small as long as the divided interval goes tiny. Based on the interval division, we have
@@
[…]
Step 3. In [18], the point-fitting function is constructed as $\tau(\theta/(\pi+n))$ based on Proposition 1. In contrast, we solve it with a discrete dynamic system that is constructed by a network made of task-driven neurons. We design a sub-network to generate a function $\phi_{2}$ mapping $k$ approximately to $f\left(\tilde{x}_{k}\right)$ for each $k$. Then $\phi_{2}(\phi_{1}(x))=\phi_{2}(k)\approx f(\tilde{x}_{k})\approx f(x)$ for any $x\in\mathcal{J}_{k}$ and $k\in\{1,2,\cdots,K\}$, which implies $\phi_{2}\circ\phi_{1}\approx f$ on $[0,1)$. $\phi_{2}$ is $T^{m(k)}(\xi)$ based on Lemma 1, where $T$ is a unimodal function from $[0,1]\to[0,1]$ that can also induce the dense trajectory.
@@
i) Given a polynomial of any order, as shown in Figure 5, through cutting, translation, flipping, and scaling, we can construct a unimodal map $T(x):[0,1]\to[0,1]$.
@@
ii) For a unimodal map $T$, we can show that for any two non-empty sets $U,V\in X$ with $\mathfrak{m}(U),\mathfrak{m}(V)>0$, there exists an $m$ such that $T^{m}(U)\cap V\neq\varnothing$. As Figure 6 shows, this is because a set $U$ with $\mathfrak{m}(U)>0$ will always contain a tiny subset $\mathcal{J}_{mi}$ such that $T^{m}(\mathcal{J}_{mi})=[0,1]$. Then, naturally $T^{m}(U)\cap V\neq\varnothing$ due to $\mathcal{J}_{mi}\subset U$.
@@
Because, given any non-empty open $U\subset X$, there is an $i\in I$ with $V_{i}\subset U$ and $k>0$ with $\xi\in f^{-k}\left(V_{i}\right)$ This means $f^{k}(\xi)\in V_{i}\subset U$. Thus the orbit of $\xi$ enters any $U$.
@@
Theorem 8 can be easily generalized into multivariate functions by using the Kolmogorov-Arnold theorem.

### v2

@@ 3.1 Overview of NS-TD
3.2 Neuronal Structure Searching
In the first stage, we aim to discover the optimal mathematical structure for the neuron’s aggregation function. Given the input vector $\mathbf{z}\in\mathbb{R}^{d}$, the function $f(\mathbf{z})$ is defined as a combination of polynomial, interaction, and periodic terms, regulated by learnable gates. The general formulation is expressed as
$f(\mathbf{z})$ $=\underbrace{\sum_{k=1}^{K}g_{k}^{(p)}\left(\mathbf{w}_{k}^{\top}\mathbf{z}^{\odot k}\right)}_{\text{Polynomial Stream}}+\underbrace{\sum_{m=2}^{M}g_{m}^{(i)}\left(\sum_{r=1}^{R}\prod_{j=1}^{m}(\mathbf{a}_{r,j}^{\top}\mathbf{z})\right)}_{\text{Interaction Stream}}+\underbrace{g^{(s)}\cdot D\sin(\mathbf{z})}_{\text{Periodic Term}},$ (2)
where $g^{(\cdot)}\in\{0,1\}$ denotes the binary gate for term selection. The first component is the Polynomial Stream, where $\mathbf{w}_{k}^{\top}\mathbf{z}^{\odot k}$ models the independent $k$-th order polynomial effect, with $\mathbf{z}^{\odot k}$ denoting the element-wise power. The second component corresponds to the Interaction Stream, which utilizes CP decomposition to approximate the $m$-th order feature interactions efficiently. Finally, the periodic term $D\sin(\mathbf{z})$ is included to capture high-frequency patterns.
3.2.1 Computation of Stream Components
In this section, we detail the computational mechanisms underlying the polynomial and interaction streams defined above. The computation for the polynomial and periodic streams is straightforward, as they primarily involve element-wise operations. The polynomial term computes $\mathbf{z}^{\odot k}$ by raising each element of the input vector to the power of $k$ independently, while the periodic term applies the sine function to the input vector. These operations maintain a linear computational complexity with respect to the input dimension.
The core computational challenge lies in the interaction stream, which models the high-order coupling between features. A naive implementation of an $m$-th order interaction would require a dense weight tensor $\mathcal{W}^{[m]}\in\mathbb{R}^{d\times\dots\times d}$ with $d^{m}$ parameters. To mitigate this exponential complexity, we employ the Canonical Polyadic (CP) decomposition. This method factorizes the high-dimensional weight tensor into a summation of rank-1 component tensors. Mathematically, the weight tensor $\mathcal{W}^{[m]}$ is approximated as
$\mathcal{W}^{[m]}=\sum_{r=1}^{R}\mathbf{a}_{r}^{(1)}\circ\mathbf{a}_{r}^{(2)}\circ\dots\circ\mathbf{a}_{r}^{(m)},$ (3)
where $R$ is the decomposition rank, $\circ$ denotes the vector outer product, and $\mathbf{a}_{r}^{(j)}\in\mathbb{R}^{d}$ represents the factor vector for the $j$-th mode of the $r$-th rank component. An illustration of a rank-$R$ CP decomposition for a third-order tensor is presented in Fig. 3.
@@ 3.2.1 Computation of Stream Components
This decomposition provides significant structural and computational advantages. By substituting Eq. (3) into the interaction term, the computation transforms from a high-cost tensor contraction into a series of efficient vector inner products, as formulated in Eq. (2). Computationally, this reduces the parameter requirements from $O(d^{m})$ to $O(mRd)$. In contrast to other decomposition methods such as Tucker decomposition, which has a complexity of $O(R^{m}+mRd)$, CP decomposition eliminates the core tensor, thereby ensuring that the model complexity scales linearly with the input dimension $d$. This efficiency allows NS-TD to handle high-order interactions in high-dimensional spaces that are otherwise computationally intractable.
3.2.2 Differentiable Structure Search
To effectively identify the optimal structural configuration from the comprehensive formulation in Eq. (2), we employ a differentiable search strategy based on the $L_{0}$ regularization framework proposed by Louizos et al. [39]. By assigning a stochastic binary gate to each candidate component in the polynomial, interaction, and periodic streams, we transform the discrete term selection problem into a continuous optimization task.
We parameterize the activation probability of each gate using a learnable structural parameter. This formulation enables the gradients to backpropagate directly from the task loss to these parameters, allowing the optimizer to jointly tune the model weights (i.e., $\mathbf{w}$ and $\mathbf{a}$) and the gate probabilities. The global objective function balances predictive accuracy with structural parsimony:
$\mathcal{J}=\sum_{(\mathbf{z},y)\in\mathcal{D}}\mathcal{L}_{task}(f(\mathbf{z}),y)+\lambda\sum_{g}\mathcal{R}(g),$ (4)
where $\mathcal{L}_{task}$ is the task-specific loss (e.g., MSE or Cross-Entropy), and the second term aggregates the expected $L_{0}$ norm penalty $\mathcal{R}(g)$ over all structural gates defined in Eq. (2). Through this mechanism, the model automatically suppresses redundant polynomial orders or interaction ranks by driving their associated gate probabilities to zero, effectively performing feature selection during training.
Crucially, to adapt this mechanism to the sensitivity of high-order polynomial and interaction learning, we implement a tailored two-phase optimization protocol. Direct application of sparsity regularization from the onset often leads to the premature pruning of complex terms (e.g., high-rank interactions) before they can capture relevant features. To mitigate this, we initially freeze the structural gates in an open state, focusing the optimization exclusively on training the model tensors to capture the dataset characterization. This warm-up phase ensures that all candidate terms, particularly those in the interaction stream utilizing CP decomposition, develop meaningful representations. In the subsequent phase, we unfreeze the gate parameters and introduce the sparsity regularization, allowing the model to prune terms based on their fully developed representational capacity.
3.3 Task-driven Neural Network Construction
Upon identifying the optimal sparse structure via the differentiable search in the first stage, we proceed to construct the task-driven neural network. This phase instantiates the discovered symbolic structure into a learnable neuronal architecture. Specifically, we extract the active index sets $\mathcal{K}$ and $\mathcal{M}$ for the retained polynomial and interaction terms, along with the inclusion status of the periodic term. Distinct from the search phase, here we assign fresh, independent trainable weights to these selected components to maximize representation learning during the final training.
Mathematically, the output $y$ of the constructed neuron is defined as
$y=\sigma\Bigg($ $\sum_{k\in\mathcal{K}}\mathbf{W}_{k}^{(p)}(\mathbf{z}^{\odot k})+\sum_{m\in\mathcal{M}}\mathbf{W}_{m}^{(i)}\left(\sum_{r=1}^{R}\prod_{j=1}^{m}(\mathbf{a}_{r,j}^{\top}\mathbf{z})\right)+\mathbf{W}^{(s)}\sin(\mathbf{z})+b\Bigg),$ (5)
where $\sigma(\cdot)$ is the activation function (e.g., ReLU). $\mathbf{W}^{(\cdot)}$ represents the re-initialized trainable weights for the preserved terms. This formulation unifies the processing of tabular and visual data: for tabular inputs, $\mathbf{W}$ indicates scalar coefficients or weight matrices; for image inputs, operations are extended to convolutions.
To ensure training stability for these higher-order architectures, we employ a hierarchical initialization strategy inspired by Chrysos et al. [8]. Parameters corresponding to linear terms (i.e., $k=1$) are initialized using a normal distribution $\mathcal{N}(0,\sigma^{2})$ to preserve variance. In contrast, weights for higher-order polynomial terms and interaction terms are initialized with values close to zero (e.g., $10^{-3}$). This approach ensures that the network initially approximates a linear model, gradually integrating non-linear and interactive complexities during the training process. The efficacy of such initialization has been validated in prior studies on quadratic neurons [15].
4 Super-expressive Property Theory of Task-driven Neurons
@@
Step 1. As Fig. 4 shows, in equal distance, divide $[0,1)$ into small intervals $\mathcal{J}_{k}=\left[\frac{k-1}{K},\frac{k}{K}\right)$ for $k\in\{1,2,\cdots,K\}$, where $K$ is the number of intervals. The left endpoint of $\mathcal{J}_{k}$ is $x_{k}$. A higher $K$ leads to a smaller approximation error. We construct a piecewise constant function $h$ to approximate $f$. The error can be arbitrarily small as long as the divided interval goes tiny. Based on the interval division, we have
@@
Step 2. As Fig. 4 shows, use the function $\phi_{1}(x)=\lfloor Kx\rfloor$ mapping all $x$ in the interval $\mathcal{J}_{k}$ to $k$, where $\lfloor\cdot\rfloor$ is the flooring function. There exists a one-to-one correspondence between $k$ and $f(\tilde{x}_{k})$. Thus, via dividing intervals and applying the floor function, the problem of approximation is simplified into a point-fitting problem. Therefore, we just need to construct a point-fitting function to map $k$ to $f(\tilde{x}_{k})$.
Step 3. In [54], the point-fitting function is constructed as $\tau(\theta/(\pi+n))$ based on Proposition 1. In contrast, we solve it with a discrete dynamic system that is constructed by a network made of task-driven neurons. We design a sub-network to generate a function $\phi_{2}$ mapping $k$ approximately to $f\left(\tilde{x}_{k}\right)$ for each $k$. Then $\phi_{2}(\phi_{1}(x))=\phi_{2}(k)\approx f(\tilde{x}_{k})\approx f(x)$ for any $x\in\mathcal{J}_{k}$ and $k\in\{1,2,\cdots,K\}$, which implies $\phi_{2}\circ\phi_{1}\approx f$ on $[0,1)$. $\phi_{2}$ is $T^{m(k)}(\xi)$ based on Lemma 1, where $T$ is a unimodal function from $0,1]\to[0,1]$ that can also induce the dense trajectory.
@@
i) Given a polynomial of any order, as shown in Fig. 5, through cutting, translation, flipping, and scaling, we can construct a unimodal map $T(x):[0,1]\to[0,1]$.
ii) For a unimodal map $T$, we can show that for any two non-empty sets $U,V\in X$ with $\mathfrak{m}(U),\mathfrak{m}(V)>0$, there exists an $m$ such that $T^{m}(U)\cap V\neq\varnothing$. As Fig. 6 shows, this is because a set $U$ with $\mathfrak{m}(U)>0$ will always contain a tiny subset $\mathcal{J}_{mi}$ such that $T^{m}(\mathcal{J}_{mi})=[0,1]$. Then, naturally $T^{m}(U)\cap V\neq\varnothing$ due to $\mathcal{J}_{mi}\subset U$.
@@
Because, given any non-empty open $U\subset X$, there is an $i\in I$ with $V_{i}\subset U$ and $k>0$ with $\xi\in f^{-k}\left(V_{i}\right)$ . This means $f^{k}(\xi)\in V_{i}\subset U$. Thus the orbit of $\xi$ enters any $U$.
@@
Theorem 6 can be easily generalized into multivariate functions by using the Kolmogorov-Arnold theorem [38].